# Explicit Visual Prompt Instruction Tuning

### v1

@@ 4.2 Main Experiments
We further evaluate our model on ViP-Bench [5], an instance-level understanding benchmark that aligns closely with the objectives of our Inst-IT. However, while ViP-Bench focuses primarily on domain-specific instance-level referring comprehension, our Inst-IT encompasses a broader scope, involving cross-instance and cross-timestamp interactions, among other aspects. As shown in Tab. 6, our model demonstrates strong generalization capabilities when evaluated on downstream tasks. Specifically, our Inst-IT with Vicuna-7B achieves comparable performance to ViP-LLaVA when using rectangular bounding boxes as visual prompts and even surpasses ViP-LLaVA when employing human-style visual prompts. Notably, our model operates as a generalist under zero-shot evaluation, whereas ViP-LLaVA benefits from in-domain tuning since it is fine-tuned on the in-domain referring comprehension dataset.
CL Tune Enc Data comb AI2D MMMU POPE GQA Inst-IT-I Next-QA VideoMME Inst-IT-V
(test) (val) (test F1) (val) (mc) (mc) (w/o subt) (mc)
All layers LLaVA-Next 61.1 35.9 86.9 61.4 45.3 56.6 45.7 31.3
All layers LLaVA-Next & Inst-IT Dataset video 60.7 34.7 86.1 61.2 60.7 59.7 47.1 43.0
All layers LLaVA-Next & Inst-IT Dataset video 62.3 35.7 86.7 62.9 61.8 62.4 46.7 44.4
None LLaVA-Next & Inst-IT Dataset video 63.1 35.0 86.9 62.5 60.2 63.2 47.2 44.3
Last 12 LLaVA-Next & Inst-IT Dataset video 63.2 34.9 87.0 62.5 60.1 63.3 47.2 44.0
Last 12 LLaVA-Next & Inst-IT Dataset (img+vid) 63.0 36.1 87.2 62.7 59.6 64.3 46.6 43.7
@@
Table 4: Ablation on data training recipe. We utilize LLaVA-Qwen2-1.5B-CLIP336 as the baseline model. Inst-IT-I and Inst-IT-V indicate the multi-choice splits of the image and video part of our Inst-IT Bench, separately.
@@
1     $+$ inst-cap & img-cap 63.0 35.1 86.1 62.7 58.9 62.4 46.0 33.8
2     $+$ temporal diff 63.0 35.6 87.1 62.7 59.6 64.2 45.6 36.9
3     $+$ video-description & qa 63.2 34.9 87.0 62.5 60.1 63.3 47.2 44.0
4     $+$ Inst-IT Dataset image 63.0 36.1 87.2 62.7 59.6 64.3 46.6 43.7
Model Synthesized visual prompts (tight bounding box) Visual prompts from human (e.g. arrow, circle)
Rec OCR Know Math Rel Lang All Rec OCR Know Math Rel Lang All
GPT-4V-turbo-detail:high [1] 58.1 69.8 59.5 71.0 61.4 51.9 60.7 56.9 69.7 63.7 80.6 61.1 45.6 59.9
GPT-4V-turbo-detail:low [1] 53.2 50.3 55.6 67.7 57.5 57.5 52.8 51.7 50.3 59.3 60.3 55.0 43.8 51.4
InstructBLIP-7B [16] 36.9 16.3 34.2 22.3 26.8 7.5 31.7 38.9 17 35.4 9.7 29.3 17.5 33.3
Shikra-7B  [9] 40.2 10.0 28.0 3.5 18.9 20.6 33.7
GPT4ROI-7B  [94] 35.6 16.7 29.7 9.7 32.5 13.8 35.1
Kosmos-2 [58] 29.5 14.2 18.5 9.7 7.5 21.9 26.9
LLaVA-1.5-7B [40] 50.8 12.4 49.2 6.5 51.8 23.8 41.6 49.1 13.0 42.9 9.7 50.0 27.5 40.2
Qwen-VL-Chat [4] 43.0 30.4 40.2 9.7 25.7 28.7 39.2 48.7 22.1 41.2 6.5 48.2 25.0 41.7
ViP-LLaVA-7B [5] 54.8 18.8 52.9 9.7 53.9 42.5 45.5 55.3 17.6 45.9 8.1 44.6 33.1 46.8
LLaVA-Next-Inst-IT-Vicuna-7B 51.3 23.7 54.2 12.9 64.3 46.2 45.1 55.0 21.3 52.5 16.1 57.5 40.6 48.2
LLaVA-Next-Inst-IT-Qwen2-7B 58.9 24.5 48.5 12.9 48.2 46.3 50.5 57.7 22.5 53.2 19.4 53.6 45.0 49.0
Table 6: ViP-Bench Evaluation Results. We direct perform evaluation with our Inst-IT models.
@@ 4.3 Ablation Study
We use Qwen2-1.5B [80] as the language model and CLIP-ViT-L-336 [61] as the vision encoder for all our ablation experiments, prioritizing efficiency. We first conduct an ablation study on the training recipe to investigate how to effectively integrate Inst-IT Dataset with existing academic SFT datasets [43] for a balanced improvement. Next, we perform a detailed analysis of the impact of each component in our Inst-IT Dataset.
Effectiveness of our continuous instruction-tuning paradigm. As shown in Tab. 4, directly mixing the video Inst-IT Dataset with LLaVA-Next-DATA leads to significant improvements on video benchmarks. However, the performance on generic image understanding slightly declines. We believe this is due to two main reasons: (1) the increased ratio of video data may suppress image understanding; (2) visually prompted images may introduce a distribution shift from natural images. To address these issues, we propose a continuous SFT paradigm based on single-image models and freeze the first 12 layers of the vision encoder to preserve realistic low-level features. As a result, our model achieves balanced performance with this training approach.
Detailed dataset combination. To further investigate the effectiveness of each component in Inst-IT Dataset, as illustrated in Fig. 3, we conduct an extensive ablation study by progressively adding data components. As shown in Tab. 5, the instance-level and image-level frame captions are essential for improving instance understanding in images. Meanwhile, temporal differences, along with video-level descriptions and QA, significantly enhance video instance understanding. Finally, incorporating the image component of Inst-IT enables our model to achieve the best-balanced performance across generic image and video understanding benchmarks, as well as our proposed instance understanding benchmark.
@@ 5 Conclusion
Instance-level understanding that detects, segments, and reasons nuanced relationships among objects has long been the goal of computer vision research, yet limited effort has been made to equip Large Multimodal Models with such capabilities. In this paper, we introduced Inst-IT Bench, a carefully curated benchmark for evaluating multimodal instance understanding abilities. Extensive evaluations for a wide range of open-source models demonstrate the limitedness of current models for understanding at the instance level. To mitigate this issue, we collected Inst-IT Dataset, the first instruction-tuning dataset with explicit instance-level visual prompting and annotation. Building upon Inst-IT Dataset, we proposed Inst-IT, a continuously fined-tuned framework that excels in instance understanding as well as widely used image and video benchmarks.
Limitation and Future Works. Our current experiments are conducted on 7B and 1.5B models due to the computation cost. Moreover, our current data pipeline is automated but constrained by the overhead of GPT-4o. We can further scale the model size and scale the dataset using a model-in-the-loop approach and improve the model through multi-round instruction tuning with self-synthesized data. We leave this direction for future work.
@@ A.1 Set-of-Marks Visual Prompting
Performing instance-level annotations is challenging, and we adopt the SoM visual prompting technique [81] to address this. Specifically, as illustrated in Fig. 4, we overlay a numeric ID at the center of each instance and maintain the same ID for a given instance across all frames. This simple augmentation can explicitly guide GPT-4o to focus more effectively on the instances of interest, enabling finer-grained and more accurate annotations. Furthermore, segmentation masks are necessary to calculate the center coordinates of each instance. Details on how these masks are obtained are provided in Sec. C.1.
@@ A.2 Prompting GPT-4o
  • The task prompt $P^{f}$ for frame-level annotation, Fig. 6.
  • The task prompt $P^{vid}$ for video-level annotation, Fig. 7.
  • The task prompt $P^{qa}$ for open-ended question-answer pairs generating, Fig. 8.
@@ B.2 LLM-based Evaluator for Open-Ended QA
Recent studies [85, 14] suggest that LLMs can serve as effective evaluators. Building on this, we use GPT-4o to assess the accuracy of open-ended question answering. Specifically, GPT-4o assigns a score between 0 and 1 based on three key factors: the question, the ground-truth answer, and the model prediction. Given that Inst-IT Bench prioritizes instance-level understanding, we pay special attention to the accuracy of instance ID references. Furthermore, for the video split of Inst-IT Bench, we emphasize the correctness of timestamps to ensure temporal correctness. The task prompt for GPT-4o is illustrated in Fig. 5.
@@ B.3 Data Example
To provide a clearer understanding of Inst-IT Bench, we present a data example in Fig. 9. Each question includes both open-ended and multiple-choice formats, focusing on specific instances or exploring the relationships and interactions between multiple instances. This design highlights the significant distinction from other benchmarks, emphasizing fine-grained understanding at the instance level.
@@ C.1 Data Collection and Processing
Collection. We select five instance segmentation datasets and two multi-object tracking datasets as sources of video data. To prevent data leakage, we only used the training splits of these datasets. Additionally, we use the SA-1B [27] dataset as a source of image data, and only utilize the first five data splits provided by officially. In total, we collect 21,418 videos and 51,101 images. Tab. 7 provides detailed statistics on the sources of our data.
@@
Processing. When constructing SoM [81] visual prompts, we need to obtain the mask annotations for each instance to determine the location of the numeric IDs. For the video instance segmentation datasets [3, 75, 59, 71, 82], the instance masks are already provided and can be used directly. For multi-object tracking datasets [35, 67], we use their bounding box annotations as prompts and generate instance masks through the SAM [27]. For images in the SA-1B dataset [27], we employ Semantic-SAM [31] to segment the instances and obtain their masks.
C.2 Data Example
@@
In this section, we provide a complete video annotation from Inst-IT Dataset to offer a clearer understanding of its content and format. In all annotations, we use the format [ID] to refer to instances and $<$timestamp$>$ to refer to timestamps. We present the frame-level annotations in Tab. 8 to Tab. 10. We can see that each frame-level annotation $Y^{f}$ consists of three parts: instance-level descriptions $y^{ins}$, image-level descriptions $y^{img}$, and temporal differences $y_{dif}$. Additionally, each video is accompanied by a series of open-ended question-answer pairs $Y^{qa}$, most of which center on specific instances or their relationships, as illustrated in Tab. 11. Furthermore, we generate a dense video-level caption $Y^{vid}$ summarizing the entire video, as shown below:
Dataset Name Ann. Type Split Sample Num.
Video Instance Segmentation
BRUST [3] mask training 500
UVO [75] mask training 5,135
OVIS [59] mask training 599
LVVIS [71] mask training 3,057
YoutubeVIS [82] mask training 2,897
Video Object Tracking
BenSMOT [35] box training 2,261
VidOR [67] box training 6,969
Image
SA-1B [59] none 1-5 51,101
Table 7: Data sources. We use 7 video datasets and 1 image dataset as our data sources. We demonstrate their annotation formats, splits we used, and the number of samples from each dataset.
Video-level caption: The video appears to document a formal or celebratory event indoors, possibly a ceremony such as a wedding or official gathering, occurring in a room with wooden paneling and art or framed pictures on the wall. At the beginning, during $<$1$>$, [1] is wearing a light gray suit and stands with [2] in a sleeveless white lace dress, and [3] in a dark floral-patterned dress. The three are close together, suggesting an intimate or focused setting. The progression between $<$2$>$ and $<$3$>$ involves subtle changes in posture and interaction. [1] moves closer to [2], appearing to hold hands or engage in an exchange, possibly involving a ring, as indicated by a bouquet of flowers. [3] remains supportive and smiling, while [5], in red, momentarily holds an object above their head, before disappearing from view by $<$4$>$.In frames $<$5$>$ to $<$7$>$, [1] and [2] maintain a close interaction, suggestive of a significant moment such as an exchange of vows or rings. They are closely observed by [3], who stands smiling nearby, while [1] and [2] occasionally adjust their positions, facing each other initially and then turning outward, which may signal transitioning from an intimate moment to posing for a photo. By $<$7$>$, [4] joins, dressed in darker attire, emphasizing the formal setting as [3] is no longer visible.Through $<$8$>$ and $<$9$>$, the group dynamics change slightly with the absence of [4] and [3] entering the scene again. [1] and [2] appear to engage in a warm interaction as [3] supports them, clapping, alongside visible hands of [4] indicating applause, marking a cheerful tone.Finally, during $<$10$>$ to $<$12$>$, the focus shifts as [1] and [2] first engage in a kiss, underscoring an intimate conclusion to their ceremony. They later stand apart slightly at the center, with [1] smiling or speaking, and [2] leaning towards [1] suggestively content. Throughout, the consistent joyous mood is accentuated by [3]’s ongoing clapping and expression of joy, emphasizing shared celebration and approval from the audience captured.”
@@
[Uncaptioned image] timestamp$<$1$>$ 1: Wearing a light gray suit with a white shirt, standing indoors. 2: Wearing a sleeveless white lace dress, holding an object in the hand. 3: Wearing a dark floral-patterned dress with long wavy hair. [1] [2] [3] are standing closely together in an indoor setting. [1] is on the left side wearing a formal, light gray suit with a white shirt. [2], in the middle, is wearing a sleeveless white lace dress, holding something in their hand. [3] is on the right side in a dark floral-patterned dress with long, wavy hair. They appear to be in a room with wooden paneling and some framed art on the wall. null
@@
[Uncaptioned image] timestamp$<$2$>$ 1: A person wearing a gray suit with a white shirt, short hair. 2: A person in a white, sleeveless dress with long dark hair. 3: A person wearing a dark floral dress with long dark hair. 5: A person wearing red, partially visible in the background. 6: A small black cellphone held in a hand. The scene appears to be in an office setting with a wooden table at the foreground. [1] is standing to the left, facing [2], and appears to be holding [2]’s finger or hand. [2] stands slightly to the right, returning focus with [1]. [3] is to the right of [2], slightly in the background, smiling and looking forward. A bouquet of white flowers lies on the table near [2]. [5] is partially visible in the background on the right, seated and wearing red. [6] is a cellphone held by [5]. Background shows a wooden wall and a reflection in a window. [1] has moved closer to [2] and is now in contact with [2]’s hand. [2] has turned slightly towards [1] compared to the previous frame. [3] remains in a similar position, but the expression suggests more engagement with the scene. [5] and [6] have appeared in the frame; [5] is visible in the background holding [6]. The table with a bouquet of flowers is now visible, indicating a shift in camera angle slightly to include more of the right side of the room.

### v2

@@ 4.3 Evaluation on other instance-understanding benchmarks
ViP-Bench [vipllava] is a region-level understanding benchmark that closely aligns with the objectives of Inst-IT. As shown in Tab.˜5, our model exhibits strong generalization performance. In particular, our Inst-IT with Vicuna-7B achieves performance comparable to ViP-LLaVA when using rectangular bounding boxes as visual prompts and even surpasses ViP-LLaVA when employing human-style visual prompts. Notably, our model performs as a generalist under zero-shot evaluation, whereas ViP-LLaVA benefits from in-domain tuning, since it is fine-tuned on the dataset of ViP-Bench.
RefCOCOg [refcocog] is a referring expression comprehension benchmark, with fewer labeling errors than its counterpart RefCOCO [refcoco]. We evaluate our LLaVA-NeXT-Inst-IT-Vicuna-7B model on this benchmark and observe a clear improvement of 10.8% over the baseline LLaVA-NeXT-Vicuna-7B (63.0% vs. 52.2%). This further confirms that our approach effectively enhances the model in instance understanding, rather than simply overfitting to our Inst-IT data format.
@@
Table 6: Ablation on data training recipe. L.N. denotes LLaVA-NeXT-Data, while Inst-IT img and Inst-IT vid refer to the image and video subsets of Inst-IT. Inst-IT-I and Inst-IT-V indicate the multi-choice splits of the image and video part of our Inst-IT Bench, respectively.
CL Tune Enc Data Combination AI2D POPE GQA Inst-IT-I Next-QA VideoMME Inst-IT-V
(test) (test F1) (val) (mc) (mc) (w/o subt) (mc)
All L.N. 61.1 86.9 61.4 45.3 56.6 45.7 31.3
All L.N. & Inst-IT vid 60.7 86.1 61.2 60.7 59.7 47.1 43.0
[…]
@@
1 ++ inst-cap & img-cap 63.0 35.1 86.1 62.7 58.9 62.4 46.0 33.8
2 ++ temporal diff 63.0 35.6 87.1 62.7 59.6 64.2 45.6 36.9
3 ++ video-description & qa 63.2 34.9 87.0 62.5 60.1 63.3 47.2 44.0
4 ++ Inst-IT Dataset img 63.0 36.1 87.2 62.7 59.6 64.3 46.6 43.7
@@ 4.4 Ablation study
We use Qwen2-1.5B [qwen2] as the language model and CLIP-ViT-L-336 [clip] as the vision encoder for ablation experiments. We first conduct ablation on the training recipe to investigate how to effectively integrate Inst-IT Dataset with existing academic SFT datasets [llavanext-data] for a balanced improvement. Next, we perform a detailed analysis of the impact of each component in our Inst-IT Dataset.
Effectiveness of our continuous instruction-tuning paradigm. As shown in Tab.˜6, directly mixing the video split of Inst-IT Dataset with LLaVA-Next-DATA leads to significant improvements on video benchmarks. However, the performance on generic image understanding slightly declines. We believe this is due to two main reasons: (1) the increased ratio of video data may suppress image understanding; (2) visually prompted images may introduce a distribution shift from natural images. To address these issues, we propose a continuous SFT paradigm based on single-image models and freeze the first 12 layers of the vision encoder to preserve realistic low-level features. Our model achieves balanced performance across both image and video benchmarks with this training approach.
Detailed dataset combination. As illustrated in figure˜2, Inst-IT Dataset contains fine-grained annotations at multi-level. To investigate the effectiveness of each component in Inst-IT Dataset, we conduct an extensive ablation by progressively adding data components. As shown in Tab.˜7, the instance-level and image-level frame captions are essential for improving instance understanding in images. Meanwhile, temporal differences, along with video-level descriptions and QA, significantly enhance video instance understanding. Finally, incorporating the image component of Inst-IT Dataset enables our model to achieve the most balanced performance across generic image and video understanding benchmarks, as well as our Inst-IT Bench.
@@ 5 Conclusion
Instance understanding that detects, segments, and reasons nuanced relationships among objects has long been the goal of computer vision research, yet limited effort has been made to equip LMMs with such capabilities. We introduced Inst-IT Bench, a carefully curated benchmark for evaluating multimodal instance understanding abilities. Extensive evaluations for a wide range of models demonstrate the limitations of current models for understanding at the instance level. To mitigate this issue, we collected Inst-IT Dataset, the first instruction-tuning dataset with explicit instance-level visual prompts and annotations. Based on Inst-IT Dataset, we proposed Inst-IT, a continuous finetuning framework that excels in instance understanding and general comprehension.
Acknowledgement This work was supported in part by the National Natural Science Foundation of China (Grant 62472098) and the Science and Technology Commission of Shanghai Municipality (No. 24511103100).
Appendix
  • In appendix˜A, we outline additional implementation details of the GPT-4o-assisted data annotation pipeline.
  • In appendix˜B, we present further information about the instance understanding benchmark, Inst-IT Bench.
  • In appendix˜C, we share more details about the instruction fine-tuning dataset, Inst-IT Dataset.
  • In appendix˜D, we provide more discussions on failure cases and real-world applications.
@@ A.1 Set-of-Marks Visual Prompting
Performing instance-level annotations is challenging, and we adopt the SoM visual prompting technique [som] to address this. Specifically, as illustrated in figure˜3, we overlay a numeric ID at the center of each instance and maintain the same ID for a given instance across all frames. This simple augmentation can explicitly guide GPT-4o to focus more effectively on the instances of interest, enabling finer-grained and more accurate annotations. Furthermore, segmentation masks are necessary to calculate the center coordinates of each instance. Details on how these masks are obtained are provided in Sec.˜C.1.
@@ A.2 Prompting GPT-4o
  • The task prompt $P^{f}$ for frame-level annotation, figure˜5.
  • The task prompt $P^{vid}$ for video-level annotation, figure˜6.
  • The task prompt $P^{qa}$ for open-ended question-answer pairs generating, figure˜7.
@@ B.2 LLM-based Evaluator for Open-Ended QA
Recent studies [mmvet, chiang2023can] suggest that LLMs can serve as effective evaluators. Building on this, we use GPT-4o to assess the accuracy of open-ended question answering. Specifically, GPT-4o assigns a score between 0 and 1 based on three key factors: the question, the ground-truth answer, and the model prediction. Given that Inst-IT Bench prioritizes instance-level understanding, we pay special attention to the accuracy of instance ID references. Furthermore, for the video split of Inst-IT Bench, we emphasize the correctness of timestamps to ensure temporal correctness. The task prompt for GPT-4o is illustrated in figure˜4.
@@ B.3 Data Example
To provide a clearer understanding of Inst-IT Bench, we present a data example in figure˜8. Each question includes both open-ended and multiple-choice formats, focusing on specific instances or exploring the relationships and interactions between multiple instances. This design highlights the significant distinction from other benchmarks, emphasizing fine-grained understanding at the instance level.
Table 8: Data sources. We use seven video datasets and one image dataset as our data sources. We show their annotation formats, the splits we used, and the number of samples from each dataset.
Dataset Name Ann. Type Split Sample Num.
Video Instance Segmentation
BRUST [brust] mask training 500
UVO [uvo] mask training 5,135
OVIS [ovis] mask training 599
LVVIS [lvvis] mask training 3,057
YoutubeVIS [ytvis] mask training 2,897
Video Object Tracking
BenSMOT [bensmot] box training 2,261
VidOR [vidor] box training 6,969
Image
SA-1B [ovis] none 1-10 51,101
@@ C.1 Data Collection and Processing
Collection. We select five instance segmentation datasets and two multi-object tracking datasets as sources of video data. To prevent data leakage, we only used the training splits of these datasets, leaving their test and validation splits untouched. Additionally, we use the SA-1B [sam] dataset as a source of image data and only utilize the first ten officially provided data splits. For each split, we only use the first 50% of its images. In total, we collect 21,418 videos and 51,101 images. Tab.˜8 provides detailed statistics on our data sources.
@@
Processing. When constructing SoM [som] visual prompts, we need to obtain the mask annotations for each instance to determine the location of the numeric IDs. For the video instance segmentation datasets [brust, uvo, ovis, lvvis, ytvis], the instance masks are already provided and can be used directly. For multi-object tracking datasets [bensmot, vidor], we prompt SAM [sam] with their bounding box annotations to generate instance masks. For images in the SA-1B dataset [sam], we employ Semantic-SAM [li2023semantic] to segment the instances and obtain their masks.
C.2 Statistics Analysis.
Number of instances. The key characteristic of Inst-IT Dataset is its specific focus on individual instances in images and videos, which provides a more fine-grained description of the visual inputs. We visualize the distribution of the number of instances in each sample in figure˜10. For the video split, each sample has an average of 3.7 instances, with a total of 79,709 instances. For the image split, each sample contains an average of 6.9 instances, totaling 351,495 instances. Across the entire dataset, each sample includes an average of 5.9 instances, adding up to 431,204 instances in total. We measure the scene complicity by the number of instances in each sample. Specifically, 31% of the samples contain $\leq$ 3 instances (simple), 39% have between 3 to 8 instances (medium), and the remaining 30% contain $\geq$ 8 instances (hard).
Dataset diversity. We visualize the object categories in Inst-IT Dataset in figure˜10, highlighting its diverse range. The objects include humans, animals, plants, vehicles, landmarks, etc. , covering domains like daily life, egocentric perspectives, sports, transportation, etc. . The rich diversity of Inst-IT Dataset ensures its applicability to real-world scenarios and enhances its transferability to different domains.
Text captions. Inst-IT Dataset contains multi-level textual descriptions of visual content, covering instances, frames, temporal changes, and video-level annotations. We conduct statistical analysis on these text annotations, including the number of each type of text, and their average length. As shown in Tab.˜9, the average length of Inst-IT Dataset is 49.1 words per caption, with video-level averaging 323.2 words, highlighting its richness of details. We also present the results of lexical analysis in Tab.˜9. The instance-level captions contain a rich variety of nouns and adjectives, indicating that they primarily describe the objects’ categories and attributes. The captions of temporal changes include a high volume of verbs and adverbs, suggesting that they capture dynamic information.
Human evaluation of data quality We invited three volunteers to rate each sample on a scale from 1 to 5, with higher scores indicating better quality. Tab.˜10 presents the scores of different types of annotations, along with the average time spent by each volunteer to evaluate each sample. The average score across all types is 4.49±0.05, indicating that the data in Inst-IT Dataset is of satisfactory quality.
Figure 9: The distribution of the number of instances per sample in Inst-IT Dataset. We separately present the distribution for the video split, image split, and the entire dataset.
Figure 10: Analysis of object categories in Inst-IT Dataset, which shows a diverse range of types spanning multiple domains.
Table 9: Statistical and lexical analysis of Inst-IT Dataset. We present the results for each annotation level as well as the entire dataset.
Caption Type #Caption #Char./Cap. #Word/Cap. #Sen./Cap. Nouns Adj. Adv. Verb. Prep.
Instance-level 836,524 102.1 24.3 1.5 26.5% 13.3% 2.3% 12.3% 10.7%
Frame-level 207,662 458.0 106.5 5.7 25.2% 10.5% 2.6% 14.9% 11.5%
Temporal-change 135,143 306.6 67.7 3.7 21.2% 10.0% 6.0% 16.4% 10.8%
Video-level 21,372 1441.8 342.2 14.3 24.8% 10.6% 3.6% 13.2% 11.8%
[…]
Table 10: Human evaluation on the quality of Inst-IT Dataset.
Instance Caption Image Caption Temporal Caption Video Caption QA Pairs
Score ($\uparrow$) 4.66±0.12 4.68±0.02 4.48±0.05 4.34±0.18 4.31±0.11
Time (s) 7.3 12.4 11.9 31.0 10.6
C.3 Data example.
@@
In this section, we provide a complete video data sample from Inst-IT Dataset to offer a clearer understanding of its content and format. In all annotations, we use the format [ID] to refer to instances and <<timestamp>> to refer to timestamps. We present the frame-level annotations in Tab.˜11. We can see that each frame-level annotation $Y^{f}$ consists of three parts: instance-level descriptions $y^{ins}$, image-level descriptions $y^{img}$, and temporal differences $y_{dif}$. Additionally, each video is accompanied by a series of open-ended question-answer pairs $Y^{qa}$, most of which center on specific instances or their relationships, as illustrated in Tab.˜12. Furthermore, we generate a dense video-level caption $Y^{vid}$ summarizing the entire video in chronological order, as shown in Tab.˜13.
Appendix D More discussions.
D.1 Failure cases.
We manually inspect the dataset and model to identify the failure cases. We find that occasional failures occur in scenarios where instances are severely occluded, the image is blurry, or instances are excessively small or crowded. These challenges are common among LMMs, and future research can further investigate them.
D.2 Real-world applications.
In real-world applications, users can interactively prompt models like SAM2 [ravi2024sam] to automatically track instances of interest and generate SoMs. Additionally, our model also supports inputs without SoMs, allowing users to specify particular instances using textual descriptions. In the first scenario, our Inst-IT introduces only a marginal overhead for generating SoMs, while in the second case, it incurs no extra cost compared to the base model.
Table 11: Inst-IT Dataset frame-level annotations. For the ease of visualization, we only demonstrate the first three frames. Please zoom in to view the instance ID labels.
@@
[Uncaptioned image] timestamp<<1>> 1: Wearing a light gray suit with a white shirt, standing indoors. 2: Wearing a sleeveless white lace dress, holding an object in the hand. 3: Wearing a dark floral-patterned dress with long wavy hair. [1] [2] [3] are standing closely together in an indoor setting. [1] is on the left side wearing a formal, light gray suit with a white shirt. [2], in the middle, is wearing a sleeveless white lace dress, holding something in their hand. [3] is on the right side in a dark floral-patterned dress with long, wavy hair. They appear to be in a room with wooden paneling and some framed art on the wall. null
@@
[Uncaptioned image] timestamp<<2>> 1: A person wearing a gray suit with a white shirt, short hair. 2: A person in a white, sleeveless dress with long dark hair. 3: A person wearing a dark floral dress with long dark hair. 5: A person wearing red, partially visible in the background. 6: A small black cellphone held in a hand. The scene appears to be in an office setting with a wooden table at the foreground. [1] is standing to the left, facing [2], and appears to be holding [2]’s finger or hand. [2] stands slightly to the right, returning focus with [1]. [3] is to the right of [2], slightly in the background, smiling and looking forward. A bouquet of white flowers lies on the table near [2]. [5] is partially visible in the background on the right, seated and wearing red. [6] is a cellphone held by [5]. Background shows a wooden wall and a reflection in a window. [1] has moved closer to [2] and is now in contact with [2]’s hand. [2] has turned slightly towards [1] compared to the previous frame. [3] remains in a similar position, but the expression suggests more engagement with the scene. [5] and [6] have appeared in the frame; [5] is visible in the background holding [6]. The table with a bouquet of flowers is now visible, indicating a shift in camera angle slightly to include more of the right side of the room.